# Towards Immersive Mixed Reality Street Play

### v2

@@ 6.1.1. Experiencing IMRSP In the Wild
A dream you dream alone is only a dream. A dream you dream together is reality.
@@ 6.2. The Frictional Path to Widespread MR Street Play Adoption
In sharp contrast, remote online observers, who lacked direct interaction and immersion, expressed far more ambivalence and skepticism towards IMRSP. Online commenters frequently voiced concerns about social awkwardness or embarrassment associated with observed behaviors, often characterizing public MR play as “strange” or “inappropriate.” Safety concerns were also prominent among these remote viewers, who worried that players might become distracted, increasing the risk of accidents or collisions in public settings. These critical perceptions resonate with earlier societal reactions to novel wearable technologies such as Google Glass, where non-users perceived the technology as intrusive or socially disruptive due to unfamiliarity and lack of direct exposure (Pelikan et al., 2024; McGrane, 1994).
@@ 6.2.2. Ethical, Safety, and Privacy Concerns in Public MR Play
Despite the positive social engagement observed during IMRSP experiences, our findings indicate significant ethical, safety, and privacy concerns (Katins et al., 2024; Millard et al., 2024) that must be addressed to facilitate broader social acceptance (Eghbali et al., 2019). A prominent issue is the asymmetric power between MR headset wearers and non-participating bystanders. MR users have privileged access to hidden digital information through outward-facing sensors, such as cameras and depth scanners, potentially recording bystanders without explicit consent or awareness. This dynamic creates a sense of vulnerability among non-users, who might perceive MR technology as intrusive or threatening their privacy—an effect reminiscent of earlier public reactions to technologies like Google Glass, where similar privacy concerns led to significant public backlash (Bajorunaite et al., 2024). Recent research also underscores the necessity for privacy-enhancing mechanisms that account for bystanders’ varying preferences for awareness and consent in everyday AR scenarios (O’Hagan et al., 2022).
@@
Furthermore, IMRSP explicitly challenges established social norms, particularly Goffman’s concept of “civil inattention,” where individuals typically ignore each other in public settings to maintain social order (Goffman, 1966). By introducing visible, unconventional behaviors into shared public spaces, MR street play inherently disrupts these existing norms, compelling bystanders to acknowledge and respond to unusual public performances. While some individuals might appreciate these playful disruptions as novel or refreshing, others could interpret them negatively, perceiving such activities as intrusive or socially inappropriate. This tension could potentially lead to social friction or resistance against MR enthusiasts, highlighting the need for proactive design and policy considerations to manage public interactions sensitively. Additional scholarship highlights that the widespread adoption of everyday AR will introduce significant ethical challenges—such as manipulation and information disorder—necessitating clear guidelines for who can mediate reality and for what purpose (O’Hagan et al., 2023). Building on this, recent work argues that everyday, pervasive AR demands a new societal consensus on “perceptual rights” to protect both users and bystanders from potential misuse and abuse (O’Hagan et al., 2024).
@@ 6.3. Methodological Implications for MR Futuring
This approach is effective because, although MR hardware will plausibly improve (e.g., smaller form factors, wider FOVs, better tracking), the core social-cognitive frictions that emerge in public, multi-party settings are likely to persist. MR technology-mediated experiences obscure what others are attending to, hindering theory of mind (McKenna et al., 2023) and making norms less visually legible to bystanders who cannot see what the wearer sees (Von Terzi et al., 2021; Desnoyers-Stewart et al., 2025). Passthrough and layered visualizations further complicate mutual understanding in public places (Bailenson et al., 2024). Additionally, MR renders reality programmable: MR designers effectively codify “strange new rules” (Rao and Protocolized, 2025)—what we call “reality protocols” (Hu, 2025)—that determine what counts as visible, actionable, or socially appropriate within an app. Because these protocols vary by app, they can drive divergent behaviors in the same physical place, further hindering theory of mind even between MR wearers. This limitation will likely persist even as hardware improves. The embodied signatures we observe—unshared referents, ambiguous gestures, and micro-breaches of street etiquette—are system-agnostic. They emerge from asymmetries of technology-mediated perception and accountability rather than from any single device’s limitations. Consequently, findings from today’s deployments speak to the long arc of MR sociality: as people encounter device-specific, app-specific rules that aren’t universally visible, they will continue to negotiate trust, accountability, and shared context in public.
@@
The future is already here—it’s just not very evenly distributed.
@@ 6.4.1. Designing for Contextual Awareness
6.4.2. Harmonizing with Social Norms through Explainable Social Affordance
@@ 6.5.2. Gameplay
Our study’s MR street play experiences were based on a single game probe—MOFA, which features three games within a magic-themed context. While this theme provided a shared framework for players to quickly understand gameplay mechanics, it may not fully capture the diverse range of potential MR street play scenarios. Although we tested only one magical fantasy theme, we believe this approach can be applied more broadly. Future studies should expand beyond magic and explore games with different themes, interaction styles, and levels of physical engagement to better understand the full spectrum of MR street play possibilities and their effects on social dynamics. We propose the following research agenda:

### v3

@@ 6.1.1. Experiencing IMRSP In the Wild
A dream you dream alone is only a dream. A dream you dream together is reality.” — Yoko Ono
@@ 6.2. The Frictional Path to Widespread MR Street Play Adoption
In sharp contrast, remote online observers, who lacked direct interaction and immersion, expressed far more ambivalence and skepticism towards IMRSP. Online commenters frequently voiced concerns about social awkwardness or embarrassment associated with observed behaviors, often characterizing public MR play as “strange” or “inappropriate.” Safety concerns were also prominent among these remote viewers, who worried that players might become distracted, increasing the risk of accidents or collisions in public settings. These critical perceptions resonate with earlier societal reactions to novel wearable technologies such as Google Glass, where non-users perceived the technology as intrusive or socially disruptive due to unfamiliarity and lack of direct exposure (Due, 2015; Denning et al., 2014; Lebeck et al., 2018).
@@ 6.2.2. Ethical, Safety, and Privacy Concerns in Public MR Play
Despite the positive social engagement observed during IMRSP experiences, our findings indicate significant ethical, safety, and privacy concerns (Katins et al., 2024; Millard et al., 2024) that must be addressed to facilitate broader social acceptance (Eghbali et al., 2019). A prominent issue is the asymmetric power between MR headset wearers and non-participating bystanders. MR users have privileged access to hidden digital information through outward-facing sensors, such as cameras and depth scanners, potentially recording bystanders without explicit consent or awareness. This dynamic creates a sense of vulnerability among non-users, who might perceive MR technology as intrusive or threatening their privacy—an effect reminiscent of earlier public reactions to technologies like Google Glass, where similar privacy concerns led to significant public backlash (Bajorunaite et al., 2024). Recent research also underscores the necessity for privacy-enhancing mechanisms that account for bystanders’ varying preferences for awareness and consent in everyday AR scenarios (O’Hagan et al., 2023a).
@@
Furthermore, IMRSP explicitly challenges established social norms, particularly Goffman’s concept of “civil inattention,” where individuals typically ignore each other in public settings to maintain social order (Goffman, 1966). By introducing visible, unconventional behaviors into shared public spaces, MR street play inherently disrupts these existing norms, compelling bystanders to acknowledge and respond to unusual public performances. While some individuals might appreciate these playful disruptions as novel or refreshing, others could interpret them negatively, perceiving such activities as intrusive or socially inappropriate. This tension could potentially lead to social friction or resistance against MR enthusiasts, highlighting the need for proactive design and policy considerations to manage public interactions sensitively. Additional scholarship highlights that the widespread adoption of everyday AR will introduce significant ethical challenges—such as manipulation and information disorder—necessitating clear guidelines for who can mediate reality and for what purpose (O’Hagan et al., 2023a). Building on this, recent work argues that everyday, pervasive AR demands a new societal consensus on “perceptual rights” to protect both users and bystanders from potential misuse and abuse (O’Hagan et al., 2024).
@@ 6.3. Methodological Implications for MR Futuring
This approach is effective because, although MR hardware will plausibly improve (e.g., smaller form factors, wider FOVs, better tracking), the core social-cognitive frictions that emerge in public, multi-party settings are likely to persist. MR technology-mediated experiences obscure what others are attending to, hindering theory of mind (McKenna et al., 2023) and making norms less visually legible to bystanders who cannot see what the wearer sees (Von Terzi et al., 2021; Desnoyers-Stewart et al., 2025; O’Hagan et al., 2023b). Passthrough and layered visualizations further complicate mutual understanding in public places (Bailenson et al., 2024). Additionally, MR renders reality programmable: MR designers effectively codify “strange new rules” (Rao and Protocolized, 2025)—what we call “(mixed) reality protocols” (Hu, 2025)—that determine what counts as visible, actionable, or socially appropriate within an app. Because these protocols vary by app, they can drive divergent behaviors in the same physical place, further hindering theory of mind even between MR wearers. This limitation will likely persist even as hardware improves. The embodied signatures we observe—unshared referents, ambiguous gestures, and micro-breaches of street etiquette—are system-agnostic. They emerge from asymmetries of technology-mediated perception and accountability rather than from any single device’s limitations. Consequently, findings from today’s deployments speak to the long arc of MR sociality: as people encounter device-specific, app-specific rules that aren’t universally visible, they will continue to negotiate trust, accountability, and shared context in public.
@@
The future is already here—it’s just not very evenly distributed.” — William Gibson
@@ 6.4.1. Designing for Contextual Awareness
6.4.2. Harmonizing with Social Norms through Explainable Social Affordances
@@ 6.5.2. Gameplay
Our study’s MR street play experiences were based on a single game probe—MOFA, which features three games within a magic-themed context. While this theme provided a shared framework for players to quickly understand gameplay mechanics, it may not fully capture the diverse range of potential MR street play scenarios. Although we tested only one magical fantasy theme, we believe this approach can be applied more broadly. Future studies should expand beyond magic and explore games with different themes, interaction styles, and levels of physical engagement to better understand the full spectrum of MR street play possibilities and their effects on social dynamics. We outline two near‑term steps: